# Sampling and estimation on manifolds using the Langevin diffusion

### v2

@@ 2.2. Langevin diffusion on a Riemannian manifold
Assume that $x\mapsto g^{ij}(x)\in C^{2,\epsilon}(M)$.
@@
No further restrictions are placed on $\phi$ (e.g., convexity). The level of smoothness imposed in these Asumptions is required primarily for proofs of Theorems 4.1-4.3, but may be weakened for the content of this section.
@@
acting on $C^{2}(M)$-functions, where $\Delta_{M}$ is the Laplace–Beltrami operator.
@@
where $B^{M}$ is a Brownian motion on the manifold $M$ defined using the Eells–Elworthy–Malliavin frame bundle construction (e.g., Hsu, 2002, p.87).
The Brownian motion $B^{M}$ can be intuited as follows: given a filtered probability space $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P}),t\geq 0$, and an $\{\mathcal{F}_{t}\}_{t>0}$-adapted $\mathbb{R}^{q}$-valued standard Brownian motion $B$, the construction uses the dynamics of $B$ on $\mathbb{R}^{q}$ to induce one on $M$ by first, mapping the dynamics of $B$ onto the orthonormal frame bundle using the notion of horizontal vector fields, and then projecting down from the frame bundle onto the manifold. Thus, in local coordinates in a chart the SDE (2.7) assumes the Itô’s form
(2.8) $\text{d}X^{i}(t)=-\frac{1}{2}g^{ij}(X(t))\frac{\partial\phi}{\partial x^{j}}(X(t))\text{d}t-\frac{1}{2}g^{kj}(X(t))\Gamma_{kj}^{i}(X(t))\text{d}t+(g^{1/2}(X(t))^{ij}(X(t))\text{d}B_{j}(t),$
@@
where $B_{j}$ are $\mathbb{R}$-valued components of $B$. Since $\phi\in C^{3,\epsilon}(M)$ and, by consequence, $g^{ij}\in C^{2,\epsilon}(M)$, (2.8) is well-defined. The frame $g^{-1/2}(x):\mathbb{R}^{q}\rightarrow T_{x}M$ maps Brownian dynamics from $\mathbb{R}^{q}$ to $T_{x}M$, and this leads to the diffusion coefficient $g^{-1/2}$ in (2.8).
@@
(2.9) $\mathrm{Ric}(v,v)+\mathrm{Hess}(v,v)^{\phi}\geq-\kappa g(v,v)\,\,\text{ for all }\,(x,v)\in TM$
for some $\kappa>0$. The condition is satisfied for any compact $M$, since the Hessian of $\phi$ is bounded under Assumption Assumption; moreover, the Ricci curvature is bounded from below since the sectional curvature at any point in $M$ is bounded from below (Bishop and Crittenden, 2011, pp. 167). We will revisit this condition in Section 5 on non-compact manifolds.
@@
Let $M=\mathbb{S}^{q}$ with canonical round metric $g=\text{d}r^{2}+\sin^{2}r\,\text{d}\theta^{2}$ with intrinsic coordinate $x=(r,\theta)$ for $r\in[0,\pi)$ and $\theta\in\mathbb{S}^{q-1}$. The Fisher–Watson distribution has the probability density function $p(x)\propto\exp(\lambda\cos^{2}r)$ with respect to the volume form for $\lambda>0$, $r\in[0,\pi)$ (Watson, 1965), and thus $\phi(x)=-\lambda\cos^{2}r$, which clearly is in $C^{\infty}(M)\subset C^{3,\epsilon}(M)$. At least, two charts are needed for the local representation of $X$ since the cut locus of every point is its antipode, and non-empty.
@@ 2.3.2. Poisson PDE
[…]
@@
Using Itô’s formula, one can derive the probabilistic representation of solution of (2.16) as (e.g., Freidlin, 1985)
@@ 3.1. Riemannian Langevin algorithm
The proposed Riemannian Langevin algorithm may be fruitfully compared to the one proposed by Grorud and Talay (1996). In their paper an Euler scheme was defined in local coordinates of a chart, and the random variables were constrained to have compact support within the chart to ensure that ${X}^{h}_{n}$ and ${X}^{h}_{n+1}$ lie in the same chart for sufficiently small $h$, to facilitate proof of weak convergence of the scheme. In contrast, our algorithm, and the subsequent proof of weak convergence, requires no such constraints (beyond the natural moment matching condition (3.1)) since the Markov chain arising from our algorithm moves directly along geodesics, and there is no need to verify that ${X}^{h}_{n}$ and ${X}^{h}_{n+1}$ belong to the same chart. In Section 6.1 we experimentally compare the use of Gaussian and discrete random variables within (3.3).
@@ 3.2. Retractions
(3.7) $\begin{bmatrix}\dot{\gamma}\\ \dot{r}\end{bmatrix}=\begin{bmatrix}r\\ -\Gamma_{ij}r^{i}r^{j}\end{bmatrix}$
@@
(3.8) $\begin{bmatrix}\gamma(0)\\ r(0)\end{bmatrix}=\begin{bmatrix}x\\ V\end{bmatrix},$
@@
$V=-\frac{1}{2}\nabla\phi(x)\sqrt{h}+g^{-1/2}(x)\xi.$
To obtain the first weak order of convergence of the Riemannian Langevin algorithm (3.4) as in the case of the exponential map (3.3), it is necessary to impose conditions on the retraction. This relates to the leading $h^{2}$ terms in the local error expansion of Lemma C.2. Consequently, we make the following assumption:
@@
To illustrate our retraction based approach, we solve the geodesic equation (3.7) by using a single step of size $\sqrt{h}$ of the standard $(1/6,1/3,1/3,1/6)$ 4th-order Runge–Kutta method (RK4) (see e.g. Hairer et al. (1993)). This is employed in Section 6.1 on the 2-sphere $\mathbb{S}^{2}$, where the exponential map is not readily available in analytic form in intrinsic coordinates $(r,\theta)$ for the round metric. Since the RK4 method approximates the solution directly in a local coordinate chart, its realisation belongs to the manifold $M$ and hence a projection onto $M$ is not required, i.e., there is no additional loss in accuracy.
@@
which is of order $O(s^{2})$ since all the terms have a derivative of $e_{1}$ and the highest derivative is of 3rd order (see (3.10)). Consequently, $D_{\dot{c}(s)}D_{\dot{c}(s)}\dot{c}(s)|_{s=0}=0$ a.s.. Thus, the RK4 approximation of the geodesic equation (2.1) satisfies Assumption Assumption.
@@ 3.3. Other Langevin-based sampling algorithms
    where $\{\eta_{n}\}$ are independent uniformly, with respect to $g$, distributed on the unit sphere $\{v\in T_{X^{h}_{n}}M:g(v,v)=1\}$ in the tangent space $T_{X^{h}_{n}}$, and $\sqrt{q}$ is used to take into account that $\mathbb{E}(\eta_{n}^{i}\eta_{n}^{j})=\delta_{ij}/q$ so that $\sqrt{q}\eta_{n}^{i}$ satisfy the conditions on moments (3.1).
@@ 4. Error bounds
Theorem 4.1 is proved for Riemannian Langevin algorithm using a retraction (3.4). For clarity of exposition, Theorems 4.2 and 4.3 are proved for the algorithm (3.3) which uses the exponential map; proofs of these can be generalised to accommodate retractions in a manner similar to the proof of Theorem 4.1.
@@ 4.1. Ensemble-averaging estimator
There are three sources of errors influencing bias of $\hat{\mu}_{\phi,N}(\varphi)$: (i) proximity to the ergodic limit controlled by a choice of $T$; (ii) the numerical integration error controlled by the time step $h$; and (iii) the Monte Carlo error controlled by the number of Monte Carlo runs $L$. The following result shows how the weak upper bound on its bias depend on both $h$ and $T$, and in a specific way on the class of functions $\varphi$. Note that the initial point $x$ in (2.8) can be random, independent of the Wiener process $w(\cdot)$, and, consequently, $X_{0}^{h}$ in the ensemble-averaging setting can be random, independent of $\xi_{n}$.
@@ 4.2. Time-averaging estimator
We remark that the error $C/T$ in (4.6)-(4.7) is due to the two sources of errors - (i) and (iii) as described before the theorem. They correspond to properties of the considered SDE (2.7). Further, it is not difficult to see that the estimates (4.6)-(4.7) imply that $\mathrm{Var}(\tilde{\mu}_{\phi,N}(\varphi))\leq C(h^{2}+1/T)$, however one can prove under a mixing condition that $\mathrm{Var}(\tilde{\mu}_{\phi,N}(\varphi))\leq C/T$ with $C>0$ independent of $h$ and $T$ (see (Mattingly et al., 2010, Section 7) for more details when $M=\mathbb{T}^{q}$).
@@ 4.4. Error extrapolation: increasing accuracy order and controlling bias
(4.9) $\lim_{N\to\infty}\mathbb{E}(\hat{\mu}_{\phi,N}(\varphi))-\mu_{\phi}(\varphi)=C_{0}h+\cdots+C_{p}h^{p}+O(h^{p+1}),$
where the constants $C_{0},\ldots,C_{p}$ are independent of $h,$ $T$, and $p\geq 0$ is an integer ($p$ can be arbitrarily large if $g$, $\phi$ and $\varphi$ belong to $C^{\infty}(M))$.
@@
(4.10) $\mu_{\phi}(\varphi)$$\approx$ $\mathbb{E}(\hat{\mu}_{\phi,N_{1}}(\varphi))+C_{0}h_{1}+O(h^{2})\,,$
$\mu_{\phi}(\varphi)$$\approx$ ${\mathbb{E}}(\hat{\mu}_{\phi,N_{2}}(\varphi))+C_{0}h_{2}+O(h^{2})\,,$
@@
(4.11) $C_{0}\approx-\frac{\mathbb{E}(\hat{\mu}_{\phi,N_{2}}(\varphi))-\mathbb{E}(\hat{\mu}_{\phi,N_{1}}(\varphi))}{h_{2}-h_{1}}+O(h)\,.$
@@
Thus, the obtained method has an accuracy of order two in $h$. We can continue exploiting (4.9) and obtain a method of order three and so on. On the other hand, $C_{0}h$ gives the leading term of the bias of $\hat{\mu}_{\phi,N_{1}}(\varphi)$, which can be used for error estimation in practice to assess the quality of the approximation.

### v3

@@ 2.2. Langevin diffusion on a Riemannian manifold
Assume that $x\mapsto g^{ij}(x)\in C^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}3},\epsilon}(M)$.
@@
No further restrictions are placed on $\phi$ (e.g., convexity). The level of smoothness imposed in these as sumptions is required primarily for proofs of Theorems 4.1-4.3, but may be weakened for the content of this section.
@@
acting on $C^{2}(M)$-functions, where $\Delta_{M}$ is the Laplace-Beltrami operator.
@@
where $B^{M}$ is a Brownian motion on the manifold $M$ defined using the Eells-Elworthy-Malliavin frame bundle construction (e.g., Hsu, 2002, p.87).
The Brownian motion $B^{M}$ can be intuited as follows: given a filtered probability space $(\Omega,\mathcal{F},\mathcal{F}_{t},\mathbb{P}),t\geq 0$, and an $\{\mathcal{F}_{t}\}_{t>0}$-adapted $\mathbb{R}^{q}$-valued standard Brownian motion $B$, the construction uses the dynamics of $B$ on $\mathbb{R}^{q}$ to induce one on $M$ by first, mapping the dynamics of $B$ onto the orthonormal frame bundle using the notion of horizontal vector fields, and then projecting down from the frame bundle onto the manifold. Thus, in local coordinates in a chart the SDE (2.7) assumes the Itô form
(2.8) $\text{d}X^{i}(t)=-\frac{1}{2}g^{ij}(X(t))\frac{\partial\phi}{\partial x^{j}}(X(t))\text{d}t-\frac{1}{2}g^{kj}(X(t))\Gamma_{kj}^{i}(X(t))\text{d}t+(g^{1/2}(X(t)){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0})}^{ij}\text{d}B_{j}(t),$
@@
where $B_{j}$ are $\mathbb{R}$-valued components of $B$. Under Assumptions Assumption and Assumption, (2.8) is well-defined. The frame $g^{-1/2}(x):\mathbb{R}^{q}\rightarrow T_{x}M$ maps Brownian dynamics from $\mathbb{R}^{q}$ to $T_{x}M$, and this leads to the diffusion coefficient $g^{-1/2}$ in (2.8).
@@
(2.9) $\mathrm{Ric}(v,v)+\mathrm{Hess}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\phi}}(v,v)\geq-\kappa g(v,v)\,\,\text{ for all }\,(x,v)\in TM$
for some $\kappa>0$. The condition is satisfied for any compact $M$, since the Hessian of $\phi$ is bounded under Assumption Assumption; moreover, the Ricci curvature is bounded from below since the sectional curvature at any point in $M$ is bounded from below (Bishop and Crittenden, 2011, pp. 167). We will revisit this condition in Section 5 on non-compact manifolds, where we also discuss how it relates to the Euclidean case.
@@
Let $M=\mathbb{S}^{q}$ with canonical round metric $g={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}r^{2}+\sin^{2}r\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}}\theta^{2}$ with intrinsic coordinate $x=(r,\theta)$ for $r\in[0,\pi)$ and $\theta\in\mathbb{S}^{q-1}$. The Fisher-Watson distribution has the probability density function $p(x)\propto\exp(\lambda\cos^{2}r)$ with respect to the volume form for $\lambda>0$, $r\in[0,\pi)$ (Watson, 1965), and thus $\phi(x)=-\lambda\cos^{2}r$, which clearly is in $C^{\infty}(M)\subset C^{3,\epsilon}(M)$. At least, two charts are needed for the local representation of $X$ since the cut locus of every point is its antipode, and non-empty.
@@ 2.3.2. Poisson PDE
It is clear that $\int_{M}(\varphi(x)-\mu_{\phi}(\varphi))\text{d}\mu_{\phi}(x)=0$, i.e., the centering condition is satisfied (e.g., Miranda, 1969; Freidlin, 1985). Under Assumptions Assumption and Assumption and $\varphi\in C^{2,\epsilon}(M)$, there is a unique (upto an additive constant) solution of the elliptic problem (2.16) which belongs to $C^{4,\epsilon}(M)$ (Aubin, 1998; Nicolaescu, 2007; Miranda, 1969) and
@@
Using Itô’s formula, one can derive the probabilistic representation of solution of (2.16) as (e.g., Freidlin, 1985; Le et al., 2024)
@@ 3.1. Riemannian Langevin algorithm
The proposed Riemannian Langevin algorithm may be fruitfully compared to the one proposed by Grorud and Talay (1996). In their paper an Euler scheme was defined in local coordinates of a chart, and the random variables were constrained to have compact support within the chart to ensure that ${X}^{h}_{n}$ and ${X}^{h}_{n+1}$ lie in the same chart for sufficiently small $h$, to facilitate proof of weak convergence of the scheme. In contrast, our algorithm, and the subsequent proof of weak convergence, requires no such constraints (beyond the natural moment matching condition (3.1)) since the Markov chain arising from our algorithm moves directly along geodesics, and there is no need to verify that ${X}^{h}_{n}$ and ${X}^{h}_{n+1}$ belong to the same chart. In Section 6.1, we experimentally compare the use of Gaussian and discrete random variables within (3.3).
@@ 3.2. Retractions
(3.7) $\begin{bmatrix}\dot{\gamma}\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\dot{v}}\end{bmatrix}=\begin{bmatrix}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v}\\ -\Gamma_{ij}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v^{i}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v^{j}}\end{bmatrix}$
@@
(3.8) $\begin{bmatrix}\gamma(0)\\ {\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v(0)}\end{bmatrix}=\begin{bmatrix}x\\ V\end{bmatrix},$
@@
$V=-\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sqrt{h}}}{2}\nabla\phi(x)+g^{-1/2}(x)\xi.$
To obtain the first weak order of convergence of the Riemannian Langevin algorithm (3.4) as in the case of the exponential map (3.3), it is necessary to impose conditions on the retraction. This relates to the leading $h^{2}$ terms in the local error expansion of Lemma B.2. Consequently, we make the following assumption .
@@
To illustrate our retraction based approach, we solve the geodesic equation (3.7) by using a single step of size $\sqrt{h}$ of the standard $(1/6,1/3,1/3,1/6)$ 4th-order Runge-Kutta method (RK4) (see e.g. Hairer et al. (1993)). This is employed in Section 6.1 on the 2-sphere $\mathbb{S}^{2}$, where the exponential map is not readily available in analytic form in the intrinsic coordinates $(r,\theta)$ for the round metric. Since this approximation occurs within a specific, localized coordinate system, the resulting solution belongs to the manifold $M$, and hence a projection onto $M$ is not required, i.e., there is no additional loss in accuracy.
@@
which is of order $O(s^{2})$ since all the terms have a derivative of $e_{1}$ and the highest derivative is of 3rd order (see (3.10)). Consequently, $D_{\dot{c}(s)}D_{\dot{c}(s)}\dot{c}(s)|_{s=0}=0$ a.s.. Thus, the approximation of the geodesic equation (2.1) by the one step of RK4 with time step $\sqrt{h}$ satisfies Assumption Assumption.
@@ 3.3. Other Langevin-based sampling algorithms
    where $\{\eta_{n}\}$ are independent uniformly distributed on the unit sphere in $\mathbb{R}^{q}$, and $\sqrt{q}$ is used to take into account that $\mathbb{E}(\eta_{n}^{i}\eta_{n}^{j})=\delta_{ij}/q$ so that $\sqrt{q}\eta_{n}^{i}$ satisfy the conditions on moments (3.1).
@@ 4. Error Bounds
Theorem 4.1 is proved for the Riemannian Langevin algorithm using a retraction (3.4). For clarity of exposition, Theorems 4.2 and 4.3 are proved for the algorithm (3.3) which uses the exponential map; proofs of these can be generalised to accommodate retractions in a manner similar to the proof of Theorem 4.1.
@@ 4.1. Ensemble-averaging estimator
There are three sources of errors of $\hat{\mu}_{\phi,N}(\varphi)$: (i) proximity to the ergodic limit controlled by a choice of $T$; (ii) the numerical integration error controlled by the time step $h$; and (iii) the Monte Carlo error controlled by the number of Monte Carlo runs $L$. The following result shows how the weak upper bound on its bias depend on both $h$ and $T$, and in a specific way on the class of functions $\varphi$. Note that the initial point $x$ in (2.8) can be random, independent of the Wiener process $w(\cdot)$, and, consequently, $X_{0}^{h}$ in the ensemble-averaging setting can be random, independent of $\xi_{n}$.
@@ 4.2. Time-averaging estimator
We remark that the error $C/T$ in (4.6)-(4.7) is due to the two sources of errors - (i) and (iii) as described before the theorem. They correspond to properties of the considered SDE (2.7). Further, it is not difficult to see that the estimates (4.6)-(4.7) imply that $\mathrm{Var}(\tilde{\mu}_{\phi,N}(\varphi))\leq C(h^{2}+1/T)$, however one can prove under a mixing condition that $\mathrm{Var}(\tilde{\mu}_{\phi,N}(\varphi))\leq C/T$ with $C>0$ independent of $h$ and $T$ (see (Mattingly et al., 2010, Section 7) for more details when $M=\mathbb{T}^{q}$). For a comparison of computational costs between ensemble averaging and time averaging see (Tretyakov, 2025, Remark 3.9).
@@ 4.4. Error extrapolation: increasing accuracy order and controlling bias
(4.9) $\lim_{N\to\infty}\mathbb{E}(\hat{\mu}_{\phi,N}(\varphi))-\mu_{\phi}(\varphi)=C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}}h+\cdots+C_{p}h^{p}+O(h^{p+1}),$
where the constants $C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}},\ldots,C_{p}$ are independent of $h,$ $T$, and $p\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2}$ is an integer ($p$ can be arbitrarily large if $g$, $\phi$ and $\varphi$ belong to $C^{\infty}(M))$.
@@
(4.10) $\mu_{\phi}(\varphi)$$\approx$ $\mathbb{E}(\hat{\mu}_{\phi,N_{1}}(\varphi))+C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}}h_{1}+O(h^{2})\,,$
$\mu_{\phi}(\varphi)$$\approx$ ${\mathbb{E}}(\hat{\mu}_{\phi,N_{2}}(\varphi))+C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}}h_{2}+O(h^{2})\,,$
@@
(4.11) $C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}}\approx-\frac{\mathbb{E}(\hat{\mu}_{\phi,N_{2}}(\varphi))-\mathbb{E}(\hat{\mu}_{\phi,N_{1}}(\varphi))}{h_{2}-h_{1}}+O(h)\,.$
@@
Thus, the obtained method has an accuracy of order two in $h$. We can continue exploiting (4.9) and obtain a method of order three and so on. On the other hand, $C_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}}h$ gives the leading term of the bias of $\hat{\mu}_{\phi,N_{1}}(\varphi)$, which can be used for error estimation in practice to assess the quality of the approximation. In Section 6.1, we apply the Talay-Tubaro extrapolation method to generate samples of the von-Mises Fisher distribution with bias of order 2.